OpenStudy (anonymous):

Randy has been awarded some money in a settlement. He has the option to take a lump sum payment of $300,000 or get paid an annuity of $2,000 per month for the next 20 years. Which is the better deal for Randy, and by how much, assuming the growth rate of the economy is 5.65% per year?

OpenStudy (anonymous):

The answer choices are Lump Sum: by $23,606.12 Lump Sum: by $16,726.56 Annuity: by $23,606.12 Annuity: by $16,726.56

OpenStudy (anonymous):

So I would be using 300000=x*1.0565^20?

OpenStudy (anonymous):

I got 99,938.1 but that does not make sens.

OpenStudy (anonymous):

sense*

OpenStudy (phi):

My thought is to compare present value of the lump sum (easy, it's 300,000) to the present value of the annuity

OpenStudy (anonymous):

Ok. So what would I be doing?

OpenStudy (anonymous):

2000((1+.0565)^(20)-1)/(.0565) I got 70,862.2

OpenStudy (phi):

Finding the present value of the annuity is like the last problem, except the annuity is paid monthly, so the rate is 5.65%/12 or 0.0565/12 and n is 20*12= 240 payments the first step is find the future value

OpenStudy (phi):

yes, right idea, but you are doing payments once a year.

OpenStudy (anonymous):

Ok. so it would be 2000((1+(.0565/12)^(20*12)-1)/(.0565/12)?

OpenStudy (phi):

yes

OpenStudy (anonymous):

So that would be 886708 so then it would be 886708-300000= 586708?

OpenStudy (phi):

886708 is the future value. But to be fair we should compare its present value to 300,000

OpenStudy (anonymous):

That is what I did right?

OpenStudy (anonymous):

Or I could compare 886708-97167.3 which is 789540.7 which is making even less sense. What am I doing wrong?

OpenStudy (phi):

where did you get 97167.3

OpenStudy (anonymous):

300000=x*(1+(.0565/12))^240

OpenStudy (phi):

300,000 is not the future value. It is a present value (how much you are paid right now)

OpenStudy (anonymous):

300000*(1+(.0565/12))^240 is 926237.41

OpenStudy (anonymous):

926237.41-886708= 39,529.41

OpenStudy (anonymous):

But that is not close to any of the answers either.

OpenStudy (phi):

yes I know. I was thinking, find the present value of the 886708, which I get to be 287197 but that leaves a difference between it and the lump sum of 300,000 - 287197= 12803 which does not match any of your choices. Have you seen a problem like this before? If so, how do they solve it ?

OpenStudy (anonymous):

I have not. This is the first time I am seeing this.. Let me look back in my textbook and see if I can see anything like it.

OpenStudy (anonymous):

The three formulas that we are given is Lump Sum PV=FV/(1+i)^nt ordinary Annuity PVOA=c((1/i)-(1/(i(1+i)^nt) annuity due PVAD= c((1/i)-(1/(i(1+i)^nt)+c

OpenStudy (phi):

I am assuming assuming the growth rate of the economy is 5.65% per year? means the interest rate is 5.65% But now I am wondering...

OpenStudy (anonymous):

I think we are supposed to use PVOA. so it would be. 2000((1/(.0565/12))-(1/((.0565/12)(1+(.0565/12))^(12*20))= 287197 so 300000-287197= 12803 But that is not an answer either.

OpenStudy (anonymous):

PVAD would be 289197 so 300000- 289197= 10803 which is not an answer wither so.

OpenStudy (phi):

The 12803 is what I got, doing it the long way....

OpenStudy (anonymous):

But that is not one of the options.

OpenStudy (anonymous):

I have class in 2 hours so I can ask her about it specifically and let her know I cannot figure it out.

OpenStudy (phi):

Does your book have any examples where they use "growth of the economy" Does it mean interest rate?

OpenStudy (anonymous):

Let me see

OpenStudy (anonymous):

There is this question "Suppose Dimitri won a "lifetime of free groceries" contest in a random drawing at his local supermarket. The specific terms of his winnings are as follows: $2,000 in groceries, by each year's end, for the next 50 years. Dimitri is thrilled with this incredible prize! But he's curious: What would the value of this prize be if he simply received one cash payment now? Assume a 4.5% annual growth rate in the economy." And they used PVOA to get 39542.02

OpenStudy (phi):

yes, the PVOA matches my original thought. It might be you are a victim of a bad question? I would show the teacher your work, and ask about the choices

OpenStudy (anonymous):

Ok. I will ask her. I have class at noon. Thank you for all your trouble!

OpenStudy (phi):

If you get an answer, please post it here, as I am curious.

OpenStudy (phi):

To recap, to compare a lump sum given immediately, to a payments stretched out over time, convert the "annuity" to Present Value (how much it is worth right now) then compare Present Value to the lump sum.

OpenStudy (anonymous):

Ok. I will post the answer when I find out! And I will ask her for help.

OpenStudy (phi):

When I use the site http://www.bankrate.com/calculators/investing/annuity-calculator.aspx with Withdrawal Amount: $2000.00 Annual Growth Rate: 5.65% Interval Between Withdrawals: Monthly Length of Annuity: 20 years Starting Principal: $288548.94 It finds 288548.94. This is 300,000-288548.94= 11451.06 less than the lump sum. This does not match any of your choices.

OpenStudy (anonymous):

My teacher thinks there is a flaw in the question. She could not get the answer they were asking for.